The marginal cost C(in dollars) to produce x bicycles is C(x)=x^2-40x+530,
A. Find the marginal cost of producing 40 bicycles

B. How many bicycles can be manufactured so that the marginal cost equals $130.
solve C(x)=130
C. Economic theory states that to maximize profit, production should continue until marginal revenue equals marginal cost. Assuming revenue equals 230, how may bicycles should be manufactured?
Thanks for help tutors

Answer by addingup(3677)

To produce 40 bikes:
A) C(40) = 40^2-40*40+530 = 530
:
B)x^2-40x+530 = 130
x^2-40x+530-530 = 130-530
x^2+(-40)x = -400 Take one half the coefficient of x (-40/2 = -20) and square it, then add to both sides:
x^2+(-40)x+400 = 0 We have a perfect square. Factor it:
(x-20)^2 = 0 Take the sqr rt on both sides
x-20 = 0
x = 20
:
You do C, I'm tired of this problem.